H2 (g) + I2 (g) ⇌ 2HI (g) 400. K for the reaction above, Kp = 62.5.


(a) If 0.40 mol of HI are introduced in 1.0 L flask heated to 400. K, calculate pressure HI (g) before any reaction.


(b) Calculate pressure ( each species when equilibrium established in (a) )


(c) H2 (g), I2 (g), HI (g) are introduced to flask at 400. K, pressures 0.40 atm, 0.40 atm, and 5.0 atm, respectively. Calculate value of Q and state direction reaction will proceed to reach equilibrium.


(d)H2 (g), I2 (g), and HI (g) are introduced to flask pressures 0.040, 0.020, 0.090 atm, respectively. If partial pressure HI (g) decreases 50%, calculate pressure of each species when equilibrium established. 


(e) For part (d) draw three curves - one for each of three gases. The curves must show how the pressure of each of the three gases changes as equilibrium is established. Label each curve with the formula of the gas. 


1
Expert's answer
2021-04-02T05:13:38-0400

Δ[NH3]=−2×Δ[N2]=−2×(0.11M)=−0.22MΔ